### v1

# TinyGraphEstimator: Adapting Lightweight Language Models for Graph Structure Inference

Abstract
Graphs provide a universal framework for representing complex relational systems, and inferring their structural properties is a core challenge in graph analysis and reasoning. While large language models have recently demonstrated emerging abilities to perform symbolic and numerical reasoning, the potential of smaller, resource-efficient models in this context remains largely unexplored. This paper investigates whether compact transformer-based language models can infer graph-theoretic parameters directly from textual graph representations. To enable systematic evaluation, we introduce the TinyGraphEstimator dataset - a balanced collection of connected graphs generated from multiple random graph models and annotated with detailed structural metadata. We evaluate several small open models on their ability to predict key graph parameters such as density, clustering, and chromatic number. Furthermore, we apply lightweight fine-tuning using the Low-Rank Adaptation (LoRA) technique, achieving consistent improvements across all evaluated metrics. The results demonstrate that small language models possess non-trivial reasoning capacity over graph-structured data and can be effectively adapted for structural inference tasks through efficient parameter tuning.
1 INTRODUCTION
Graphs are a fundamental representation for modeling complex relationships in domains such as social networks, biology, and information systems. Accurately characterizing these networks requires estimating key graph-theoretic properties - such as connectivity, clustering, and chromatic number - that capture their structure and complexity. Traditional graph-learning approaches, including Graph Neural Networks and algorithmic estimators, achieve strong results but depend on task-specific architectures and explicit structural features.
To address this gap, we investigate whether small transformer-based language models can infer structural graph parameters directly from serialized graph representations. The models operate on deterministic graph formats - edge lists or adjacency lists - and are tasked with predicting graph-theoretic measures from these structured inputs. This setup allows us to examine whether compact language models can generalize reasoning patterns over relational structures without relying on large-scale model capacity or specialized graph algorithms.
We evaluate several small, resource-efficient models on this task and further adapt them using the Low-Rank Adaptation (LoRA) fine-tuning method. The results show that while the models exhibit limited capability in the zero-shot setting, fine-tuning leads to consistent and substantial improvements across all structural measures. These findings demonstrate that compact language models, when properly adapted, can acquire reasoning skills for graph-structured data while maintaining high computational efficiency.
The main contributions of this paper are as follows:
  • We present a systematic evaluation of small, open language models on the task of inferring graph-theoretic parameters from structured graph representations.
  • We demonstrate that LoRA-based fine-tuning yields consistent performance improvements, highlighting the potential of compact models for graph analysis tasks.
  • We introduce the TinyGraphEstimator dataset, a balanced and publicly available collection of connected graphs annotated with key structural properties for reasoning evaluation.
2 RELATED WORK
The task of predicting graph-theoretic properties has traditionally been addressed using specialized graph learning architectures such as Graph Neural Networks (GNNs) [Kipf and Welling, 2016, Hamilton et al., 2017, Wu et al., 2019]. These models operate directly on adjacency structures to learn node embeddings and global representations that capture connectivity and topology. Numerous studies have explored the prediction of structural metrics including clustering coefficients, path lengths, and degree distributions from learned embeddings [You et al., 2018, Dwivedi et al., 2022]. However, while GNNs provide strong inductive biases for graph tasks, they require task-specific supervision and often lack general reasoning ability outside the training distribution. Our work diverges from these approaches by treating the inference of graph properties as a language-based reasoning problem, where information is expressed textually and interpreted by small language models.
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [Brown et al., 2020, Bubeck et al., 2023]. Studies have shown that transformer-based architectures can internalize relational structures [Wang et al., 2025] and perform graph-related reasoning when provided with suitable prompts or structured inputs. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 5B parameters - can infer precise graph-theoretic quantities such as node count, triangle number, or chromatic number. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
A growing body of work explores how language models can represent or reason about graphs through textual or hybrid encodings. GraphGPT [Tang et al., 2024] and related studies encode adjacency information as token sequences, enabling LLMs to reconstruct or classify graph structures. Other works have investigated relational reasoning benchmarks [Hu et al., 2025, Agrawal et al., 2024], showing that LLMs can approximate certain structural statistics with appropriate context. Our approach builds on this foundation but emphasizes the systematic evaluation of structural parameter inference - quantitative graph descriptors such as density, transitivity, and global efficiency - using small models specifically adapted for this purpose.
Fine-tuning language models for specialized reasoning tasks poses significant computational and memory challenges. Parameter-efficient adaptation techniques such as Low-Rank Adaptation (LoRA) [Hu et al., 2021], prefix tuning [Li and Liang, 2021], and adapter-based methods [Houlsby et al., 2019] have emerged as scalable alternatives that modify only a small subset of parameters while preserving base model capabilities. These techniques have been successfully applied across a range of domains including code understanding [Han et al., 2024], reasoning [Bi et al., 2024], and domain adaptation [Lu et al., 2024]. In our study, we employ LoRA to specialize small language models for structural graph inference, demonstrating consistent performance improvements across all measured graph parameters. This confirms that lightweight fine-tuning can effectively enhance reasoning abilities even in compact architectures.
In contrast to prior work that primarily investigates either graph representation learning or large-scale emergent reasoning, our research unifies both directions by systematically testing the graph reasoning capacity of small, adaptable language models. By combining structured graph representations with efficient fine-tuning, our approach provides new insights into how compact LLMs internalize relational structure and approximate graph-theoretic computations.
3 PROPOSED SOLUTION
The proposed approach evaluates and enhances the ability of small language models to infer a wide range of graph-theoretic parameters directly from structured graph representations. Each model is prompted to predict a structured set of key structural properties that collectively describe its topology and complexity, based on the graph’s edge list or adjacency list representation. The inferred parameters include the number of nodes, edges, and density, as well as detailed degree-based statistics such as minimum degree, mean degree, maximum degree, and degree standard deviation. In addition to these fundamental attributes, we evaluate the models’ ability to estimate higher-order structural metrics including the total number of triangles, average clustering coefficient, transitivity, average shortest path length, diameter, chromatic number, and global efficiency. Collectively, these parameters capture both local and global aspects of graph organization, providing a comprehensive benchmark for assessing reasoning over graph structures.
In the first stage of our study, we test unmodified (plain) small language models to evaluate their intrinsic ability to perform numerical reasoning and pattern recognition over graph-structured data. This zero-shot evaluation establishes a baseline for each model’s inherent capacity to approximate structural graph measures from textual inputs. In the second stage, we apply lightweight fine-tuning using the Low-Rank Adaptation (LoRA) method, which enables efficient parameter adjustment without retraining the full model. This adaptation significantly improves performance across all evaluated parameters, demonstrating that even compact models can be effectively specialized for graph inference through minimal, targeted fine-tuning. The resulting framework provides a scalable and efficient solution for structural property estimation of graphs, highlighting the latent reasoning capabilities of small language models when guided by domain-specific adaptation.
3.1 Models Selection
To investigate the ability of small language models to infer structural properties of graphs, we selected a group of compact yet high-performing models that combine efficient deployment with strong reasoning capacity. Our goal was to identify models that are powerful enough to capture graph-theoretic patterns while remaining computationally accessible for systematic experimentation. Model selection was guided by comparative performance metrics reported on the LLM-Stats leaderboard [LLM Stats, 2024], focusing on models with fewer than 5 billion parameters that demonstrate competitive reasoning accuracy across diverse benchmarks. In particular, we prioritized models offering extended context lengths of at least 128,000 tokens, a crucial requirement for processing full graph representations.
The final selection includes Qwen-2.5-3B [Yang et al., 2025], Llama 3.2-3B [Grattafiori et al., 2024], and Phi-4-mini-4B [Abdin et al., 2024] (all in instruct variant). These models were chosen because they consistently rank among the strongest sub-5B architectures in reasoning and comprehension tasks while maintaining full open availability for reproducible research. Their extended context capacity allows complete representations of graphs to be processed in a single inference window. Moreover, their compact architecture enables efficient fine-tuning and evaluation on standard hardware, making them ideal candidates for exploring how small-scale language models internalize and generalize graph-structural information.
3.2 Dataset Construction
To evaluate the ability of small language models to infer graph-theoretic properties, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [Erdös and Rényi, 1959]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [Barabási and Albert, 1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [Watts and Strogatz, 1998].
For reproducibility, the following parameter ranges were used during dataset generation. In the ER model, the edge probability $p$ was sampled from the interval $\log(n)/n+0.01,0.35]$. In the BA model, the number of edges to attach from a new node $m$ was sampled from $1,\min(6,n-1)]$. In the WS model, the even neighborhood size $k$ was drawn from $2,\min(n-2,12)]$ and the rewiring probability $\beta$ from $[0.05,0.35]$. Each model was sampled uniformly across its respective parameter range, producing connected, non-degenerate graphs suitable for evaluating structural inference capabilities.
For each generated instance, the number of nodes $n$ is uniformly sampled from the range $[20,30]$, and connectivity is enforced by unbiased resampling until a connected instance is obtained, avoiding bias introduced by artificial augmentation. Each model contributes an equal share of graphs, ensuring diversity in structural characteristics such as degree distribution, clustering, and connectivity patterns.
Graphs are stored in multiple formats (edgelist, adjacency list, and JSON node-link representation), accompanied by both compact and full metadata that capture key structural properties including node and edge counts, density, degree statistics, clustering coefficient, transitivity, chromatic number, diameter, and triangle count. The dataset consists of 1,000 graphs for training and 100 graphs for testing, evenly distributed across the four models. This design provides a well-controlled yet structurally varied benchmark for assessing the generalization and reasoning abilities of lightweight language models in inferring global and local graph parameters.
All generated graphs, along with their metadata and manifest files, constitute the TinyGraphEstimator dataset, which is one of the key contributions of this work. It is publicly available to support reproducibility and further research on graph-structured reasoning in language models. The dataset provides a balanced and well-documented collection of connected graphs spanning diverse generative models, enabling systematic evaluation of inference performance across structural regimes.
3.3 Training Setup
We fine-tune each base model with parameter-efficient supervised instruction tuning using the TRL SFTTrainer [von Werra et al., 2020] and LoRA adapters. Training data consist of paired promptcompletion examples: the prompt encodes the normalized graph (edge or adjacency list), the fixed target schema, and concise instructions; the completion is the gold JSON with exactly fourteen fields. To avoid learning spurious prompt tokens and to sharpen output formatting, we use completion-only loss. LoRA adapters are applied to attention and MLP projection modules with rank r=16r{=}16, $\alpha{=}32$, and dropout $0.05$, leaving the base weights frozen. We train with sequence length $2048$, batch size $2$ and gradient accumulation $8$, cosine learning-rate schedule with warmup ratio $0.03$, learning rate $2\!\times\!10^{-4}$, and one epoch over the training set. Mixed precision uses bfloat16 when available.
Compute Configuration
All experiments (training and validation) were executed on a single NVIDIA RTX 3090 GPU (Ampere, 24 GB GDDR6X, Tensor Cores) in a standard Linux x86_64 environment.
3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata, detects whether the input is an edge list or adjacency list, and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (nodes, edges, density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with optional JSON-constrained decoding via lm-format-enforcer [Gat, 2024] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema (type enforcement with graceful acceptance of null when a value is not produced). Accuracy is computed per field: integers require exact equality (after integer casting), while real-valued fields use absolute and relative tolerances to decide correctness and record signed errors.
Table 1: Normalized Root Mean Squared Error (NRMSE) for graph parameter inference across small language models. Each model is evaluated in plain (zero-shot) and fine-tuned variants. Values are normalized by the range of each parameter and averaged over the test set.
Parameter Llama-3.2-3B Qwen-2.5-3B Phi-4-mini-4B
Plain Fine-tuned Plain Fine-tuned Plain Fine-tuned
Nodes 2.36 0.51 2.36 0.37 2.36 0.38
Edges 0.54 0.18 0.54 0.33 0.54 0.27
Density 0.69 0.24 0.58 0.22 0.64 0.32
Degree Min 0.43 0.23 0.43 0.35 0.43 0.23
Degree Mean 0.63 0.22 0.63 0.32 0.63 0.33
Degree Max 0.58 0.26 0.58 0.19 0.58 0.35
Degree Std 0.52 0.37 0.54 0.27 0.54 0.41
Triangles Total 0.32 0.23 0.32 0.29 0.32 0.30
Average Clustering 0.54 0.20 0.54 0.19 0.54 0.23
Transitivity 0.54 0.24 0.55 0.23 0.55 0.24
Average Shortest Path Length 0.35 0.14 0.35 0.09 0.41 0.23
Diameter 0.28 0.13 0.29 0.08 0.29 0.22
Chromatic Number 0.69 0.19 0.55 0.25 0.70 0.30
Global Efficiency 1.01 0.21 1.37 0.16 1.06 0.31
4 RESULTS
The experimental results, summarized in Table 1, highlight consistent trends in the ability of small language models to infer graph-theoretic properties. In the zero-shot setting, all models perform poorly across most parameters, indicating limited inherent capability for structural reasoning. While simple counts are occasionally inferred correctly, higher-order statistics - particularly those involving multi-node dependencies or global structure, such as triangles, clustering coefficients, path lengths, and chromatic number - show substantial deviations from the ground truth. This observation confirms that, without explicit adaptation, compact models lack the internal procedural knowledge required for accurate structural reasoning. Quantitatively, both the mean absolute error (MAE) and root mean squared error (RMSE) are high across almost all metrics, reflecting limited numerical precision and inconsistency between samples.
After parameter-efficient fine-tuning using the LoRA approach, performance improves consistently for every evaluated field. The reductions in MAE and RMSE across model families demonstrate that even lightweight adaptation can significantly enhance reasoning over graph-structured data. Improvements are most pronounced for mid-level structural descriptors - degree statistics, density, and clustering-related metrics - indicating that fine-tuning enables models to internalize relational patterns and approximate algorithmic reasoning without direct supervision over computation steps. More complex properties such as average shortest path length and chromatic number remain challenging but also show measurable gains. Overall, the adapted models produce more stable, schema-consistent, and numerically coherent predictions, suggesting that small transformers can, when properly guided, develop an emergent understanding of graph topology sufficient for practical inference tasks. These results collectively validate the effectiveness of the TinyGraphEstimator dataset and the LoRA fine-tuning strategy in unlocking latent reasoning capacity within compact language models.
5 LIMITATIONS
While this study demonstrates the potential of small language models for graph-structural reasoning, several limitations should be acknowledged. First, although LoRA fine-tuning consistently improves inference accuracy, the models still exhibit challenges in predicting complex or global parameters such as chromatic number and average shortest path length, which require deeper combinatorial reasoning. Second, the evaluation is based on synthetic graphs generated from a selected set of well-established random graph models, which provide controlled variability but may not reflect the full diversity of real-world network structures. Incorporating additional graph types or domain-specific datasets in future work could further strengthen the assessment of model generalization. Third, the experiments rely on serialized graph formats - edge lists or adjacency lists - that, while simple and interpretable, may not capture complex or densely connected structures as effectively as more expressive representations. Future work could explore alternative input forms or hybrid architectures that integrate symbolic and neural reasoning. Despite these limitations, the results highlight promising directions for lightweight, adaptable language models in graph-theoretic inference.
6 CONCLUSION
This work explored the capability of small language models to infer a wide range of graph-theoretic parameters directly from structured graph representations. Using the proposed TinyGraphEstimator dataset, we systematically evaluated compact transformer-based models on tasks involving both local and global structural reasoning. The results demonstrate that even models with fewer than 5 billion parameters possess capacity to approximate structural graph measures when adapted through parameter-efficient fine-tuning using the LoRA method. Across all evaluated metrics, fine-tuned models consistently outperformed their zero-shot baselines, confirming that lightweight adaptation can effectively enhance structured reasoning without the computational overhead of large-scale models.
Overall, this study provides new insight into how small language models can internalize and generalize graph-structural information. By combining efficiency, adaptability, and interpretability, such models offer a promising foundation for scalable graph reasoning and for extending language-based inference to structured, symbolic, and relational domains.
LLM USAGE STATEMENT
This manuscript acknowledges the use of ChatGPT [OpenAI, 2022], powered by the GPT-5 language model developed by OpenAI, to improve language clarity, refine sentence structure, and enhance overall writing precision.